Due to a failing power supply we lost half of our switching capacity as well as a power feed in one our server racks.
This might have caused some temporarily higher latency and/or packetloss between 00:00 - 03:00h CET for a few HDD/NVME VPS nodes.
A workaround has been applied to restore full switching capacity and permanent fixes are currently being made.